OpenStudy (anonymous):

A ball is thrown straight down from the top of a 300-ft building with an initial velocity of -12ft per second. The position function is s(t)=-16t^2+v0t+s0. What is the velocity of the ball after 4 seconds?

OpenStudy (paxpolaris):

\[s(t)=-16t^2+v_0t+s_0\] v0= initial velocity = -12 s0 initial height = 300

OpenStudy (anonymous):

answer would be -140?

OpenStudy (anonymous):

????????

OpenStudy (paxpolaris):

\[s(t)=-16t^2-12t+300\]

OpenStudy (paxpolaris):

plug in 4 \for t and i am getting negative 4 as the answer

OpenStudy (anonymous):

s'(t)=-32t-12 = -32(4)-12 =-140

OpenStudy (paxpolaris):

sorry velocity ... yes -140 .